# Lite-Mind: Towards Efficient and Versatile Brain Representation Network

### 06 Dec 2023 v1

@@ Abstract
Research in decoding visual information from the brain, particularly through the non-invasive fMRI method, is rapidly progressing. The challenge arises from the limited data availability and the low signal-to-noise ratio of fMRI signals, leading to a low-precision task of fMRI-to-image retrieval. State-of-the-art MindEye remarkably improves fMRI-to-image retrieval performance by leveraging a deep MLP with a high parameter count orders of magnitude, i.e., a 996M MLP Backbone per subject, to align fMRI embeddings to the final hidden layer of CLIP’s vision transformer. However, significant individual variations exist among subjects, even within identical experimental setups, mandating the training of subject-specific models. The substantial parameters pose significant challenges in deploying fMRI decoding on practical devices, especially with the necessitating of specific models for each subject. To this end, we propose Lite-Mind, a lightweight, efficient, and versatile brain representation network based on discrete Fourier transform, that efficiently aligns fMRI voxels to fine-grained information of CLIP. Our experiments demonstrate that Lite-Mind achieves an impressive 94.3% fMRI-to-image retrieval accuracy on the NSD dataset for Subject 1, with 98.7% fewer parameters than MindEye. Lite-Mind is also proven to be able to be migrated to smaller brain datasets and establishes a new state-of-the-art for zero-shot classification on the GOD dataset. Code is available at https://github.com/gongzix/Lite-Mind.
@@ 1 Introduction
Figure 1: The significance of lightweight brain decoding model. Clouds of different colors represent different subject-wise models, and the size of the cloud represents the size of the model. The storage cost of MindEye is extremely high, and each person’s model parameter quantity is at the GB level, while Lite-Mind only requires the MB level parameter quantity.
The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images. However, fMRI data often suffers from spatial redundancy, noise, and sample sparsity, leading to poor representation of fMRI signals and potential overfitting to noise distribution for deep models. To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models [27, 15, 9, 1, 13, 44]. Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models [28] as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures [21, 23, 19, 33]. Initially, the CLS token of the CLIP image encoder was used to align voxels in Mind-Reader [19]. Considering the CLS token alone provides limited information for fMRI data, BrainClip [21] and UniBrain [23] introduce the text annotations of images to assist in aligning fMRI voxels with image representations. Brain-diffuser [25] aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
@@
To tackle the trade-off between model capacity and scalability in fMRI-to-image retrieval, we propose Lite-Mind, a lightweight, efficient, and versatile brain representation network. We redesign MindEye with our elaborate Discrete Fourier Transform (DFT) Backbone, avoiding the heavy MLP Backbone used in MindEye. Extensive experiments show that Lite-Mind achieves 94.3% retrieval accuracy for Subject 1 on the NSD dataset, with 98.7% fewer parameters than MindEye. Lite-Mind also proves its adaptability to smaller brain datasets and establishes a new state-of-the-art zero-shot classification on the GOD dataset. Our contributions are summarized as follows:
  • We empirically show that it is unnecessary to overly focus on each voxel value of fMRI-to-image retrieval while focusing on the global features of fMRI patches is more lightweight and efficient.
  • We propose a new Brain Representation Network DFT Bcakbone based on Discrete Fourier Transform, as an efficient and versatile fMRI representation network for fMRI-to-image retrieval. Its effectiveness, efficiency, and lightweight are demonstrated.
  • We migrated DFT Backbone to a smaller GOD dataset and demonstrated its universality. It achieves state-of-the-art zero-shot classification performance.
@@ 2.1 Brain Visual Decoding
The study of visual decoding using fMRI in the human brain has been a long-standing endeavor. In the most important fMRI-image reconstruction and partial fMRI-image retrieval tasks, the foundation of performance is how to align fMRI signals to the image representation space or intermediate space. Due to the low signal-to-noise ratio of fMRI, the initial research emphasized using models based on linear regression to extract fMRI signals and images into the intermediate space for functional decoding of the human brain or reconstruction of faces and natural scenes. Further research utilizes pre-trained VGG to enrich hierarchical image features [13, 34, 7]. With the development of cross-modal tasks, the image-text pre-training model CLIP has been introduced into fMRI-image research. Mind-reader [19] is the first to adopt a contrastive learning approach, using the nsdgenal ROI region of NSD data and a shallow resnet-like model to align fMRI with the CLS token output by the CLIP image encoder. Mind-Reader did not achieve good performance in fMRI-image retrieval when using InfoNCE loss. BrainClip [21] uses a VAE model for retrieval tasks while still using the CLS token. Brain diffuser [25] introduced a fine-grained representation of the image of the last hidden layer in CLIP. But they used 257 different linear regression models to align each layer of the 257 hidden layers, resulting in the model being too cumbersome. Mind-Vis [5] conducted self-supervised mask pre-training on fMRI signals for another dataset, BOLD5000, and demonstrated that voxel patches can be used to process fMRI voxel values. Using a linear regression model, Takagi et al. [36] introduce additional ventral brain regions and align fMRI to the Stable Diffusion variable space. MindEye [33] uses a deep MLP backbone for contrastive learning, aligning the last hidden layer $257\times 768$ of CLIP, and using diffusion prior in DALLE·2 [29] for the first time to narrow the disjointed vectors. At the same time, it creatively proposed a mixed contrast loss MixCo as a data augmentation method to make the large model converge. However, to the best of our knowledge, no previous research has focused on lightweight and efficient brain networks.
@@ 2.2 Fourier Transform in Deep Learning
Fourier transform plays a vital role in the area of digital signal processing. It has been introduced to deep learning for enhanced learning performance [11, 8, 43]. GFNet [30] utilizes fast Fourier transform to convert images to the frequency domain and exchange global information between learnable filters. As a continuous global convolution independent of input resolution, Guibas et al. [12] design the adaptive Fourier neural operator frame token mixing. Xu et al. [41] devise a learning-based frequency selection method to identify trivial frequency components and improve the accuracy of classifying images. On text classification, Lee-Thorp et al. [18] use the Fourier transform as a text token mixing mechanism. Furthermore, the Fourier transform is also applied to forecast time series  [3, 17, 16, 42]. To increase the accuracy of multivariate time-series forecasting, Cao et al. [3] propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the spectral domain. Yang et al. [42] propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum. Wang et al. [38] prove that frequency domain MLP is more efficient than real domain MLP. fMRI is essentially a signal that can be processed in the frequency domain, but to the best of our knowledge, there has been no research on applying FFT to efficient brain representation.
3 Preliminaries
3.1 Motivation and Analysis
FMRI signals are considered to have spatial redundancy, noise, and sample sparsity; even the nsdgeneral of NSD data has been preprocessed, which makes huge fully connected MLP of MindEye consume a lot of unnecessary parameter quantities. Furthermore, various data augmentation methods in MindEye are only used to prevent the overfitting of large models, which is also caused by MLP’s excessive focus on each voxel value. The fine-grained features come from the rich representation of the image by the last hidden layer of CLIP. Aligning multiple representations does not require excessive attention to each voxel’s values but rather to the different patterns of the entire brain region.
The diffusion model that approximates disjointed cross-modal vectors is effective for retrieval. This paper does not discuss the lightweight of the diffusion model but rather discusses the efficient brain representation network backbone. We focus on designing a lightweight voxel encoder that can quickly capture global information of fMRI and efficiently align fMRI voxels into fine-gained image representations.
@@ 4 Lite-Mind
Figure 2: Overview of our proposed Lite-Mind. Figure (a) is the architecture of the MLP Backbone of MindEye with 4 residual blocks(4096×4096). Figure (b) is our architecture of the DFT backbone and pipeline of retrieval. FMRI voxels are represented through DFT Backbone to obtain voxel embeddings. On the upper right is the overall retrieval pipeline, where LAION-5B retrieval and zero-shot classification are based on CLS token and diffusion projector. At the bottom right is the filter pyramid aggregation method, reducing the receptive field to align CLIP representation.
@@ 4.1 Overview
As illustrated in Figure 2, Lite-Mind comprises three components: 1) DFT Backbone maps flattened voxels to an intermediate space with a dimension of 257 × 768, corresponding to the last hidden layer of CLIP ViT/L-14. 2) Retrieval pipeline contains various downstream retrieval-based tasks with the voxel embeddings from the DFT backbone. 3) Contrastive Learning aims to bring the representation of two modalities closer together.
@@ 4.2 DFT Backbone
The backbone consists of patch partitioning followed by multilayer filter blocks, while the filter block adopts a pyramid aggregation strategy and finally outputs voxel embeddings. Given a fMRI-image sample pair [$x_{i}$,$y_{i}$] ($x_{i}$ represents the $i$-th fMRI sample, $y_{i}$ represents the image paired with it). Our optimization objectives are as follows:
@@
Patchfy. Given a fMRI-image pair $x_{i}=[v_{1},v_{2},...,v_{m}]$ ($i$ represents the $i$-th fMRI sample, $m$ is the length of fMRI voxels). we firstly divide $x_{i}$ into $n$ non-overlapping region patches $p_{i}=[p_{1},p_{2},...,p_{n}](m=nl)$ and adopt CNN to generate the multi-channel representation of voxel patches:
$p^{c}_{i}=PE(CNN(x_{i}))$ (2)
Where $PE(\cdot)$ is the region patch embedding for voxels, and $c$ is the number of patch channels.
Filter Block. After we obtain the voxel patch embeddings $x_{i}=[p^{c}_{1},p^{c}_{2},...,p^{c}_{n}]$, then the spectrum of voxel patch embeddings is obtained by multi-channel 1D DFT as below:
$X^{c}[k]=F_{pat}(x[i])=\sum_{i=0}^{n-1}x[i]e^{-j(2\pi/m)ki}$ (3)
Where $X\in\mathbb{C}^{n\times d}$ is a complex tensor, $F_{pat}(\cdot)$ denotes the 1D DFT along the voxel patch dimension. Note that self-attention computes the spatial dependencies in a quadratic time complexity, while DFT can be efficiently implemented via a fast Fourier transform in logarithmic time complexity.
After obtaining the spectrum of voxel patches, we use a voxel filter library for filtering processing. Voxel spatial features are effectively consolidated within each frequency element, enabling the extraction of informative features from voxels through the point-wise product in the frequency domain. To expand voxel vectors to final hidden layers of CLIP with dimensions much larger, we introduce a filter library for each channel of voxel patches $X^{c}$, to obtain fine-grained alignment of images in the frequency domain. We use $K=[K_{c_{1}},K_{c_{2}},...,K_{c}]$ to represent the filter library, where c is the number of filters in the filter library:
$\hat{X}=\sum_{i=1}^{c}\frac{1}{l}X^{i^{2}}\odot k_{i}cos(\frac{(2i-1)\pi}{2k})$ (4)
where $\odot$ is the element-wise multiplication, $|X^{i}|^{2}$ is the power spectrum of $X^{i}$, $l$ is the length of $X^{i}$. The $|X^{i}|^{2}$ operation smooths the spectrum, highlighting the main components of the fMRI spectrum from a global perspective. It also facilitates voxel information concentration and noise filtering. $cos((2i-1)\pi/2k)$ compacts better energy and can aggregate the more important information in fMRI voxels. Its combination of application with the channel filter library $K$ allows for various fine-grained features alignment to images from different channel modes.
Filter Pyramid Merge. To better align the voxel patches with the fine-grained representation of the image in the frequency domain, we adopt a filter pyramid merging strategy to expand the receptive field of the voxel patch channels. Gradually attenuated filter library converges to obtain final hidden layer $257\times 768$ for alignment, as shown in the lower right corner of Figure 2.
4.3 Retrieval pipeline

### 12 Mar 2024 v2

# Lite-Mind: Towards Efficient and Robust Brain Representation Learning

@@ Abstract
Research in decoding visual information from the brain, particularly through the non-invasive fMRI method, is rapidly progressing. The challenge arises from the limited data availability and the low signal-to-noise ratio of fMRI signals, leading to a low-precision task of fMRI-to-image retrieval. State-of-the-art MindEye remarkably improves fMRI-to-image retrieval performance by leveraging a large model, i.e., a 996M MLP Backbone per subject, to align fMRI embeddings to the final hidden layer of CLIP’s Vision Transformer(ViT). However, significant individual variations exist among subjects, even within identical experimental setups, mandating the training of large subject-specific models. The substantial parameters pose significant challenges in deploying fMRI decoding on practical devices. To this end, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm based on Discrete Fourier Transform, that efficiently aligns fMRI voxels to fine-grained information of CLIP. Our experiments demonstrate that Lite-Mind achieves an impressive 94.6% fMRI-to-image retrieval accuracy on the NSD dataset for Subject 1, with 98.7% fewer parameters than MindEye. Lite-Mind is also proven to be able to be migrated to smaller fMRI datasets and establishes a new state-of-the-art for zero-shot classification on the GOD dataset. Code will be available at https://github.com/gongzix/Lite-Mind.
@@ 1 Introduction
Figure 1: The significance of lightweight brain decoding model. Subject-specific models are visually represented by clouds of distinct colors, and the cloud size correlates with model parameters.
The success of fMRI-to-image retrieval heavily relies on aligning fMRI signals with the representation space of images. However, fMRI data often suffers from spatial redundancy, noise, and sample sparsity, leading to poor representation of fMRI signals and potential overfitting to noise distribution for deep models. To address these challenges, previous studies have found ridge regression and shallow linear models to be effective solutions for fMRI-based models [29, 15, 9, 1, 13, 48]. Recently, it has been demonstrated the effectiveness of leveraging pretrained CLIP models [30] as a powerful bridge between fMRI voxels and images, since CLIP’s image embeddings capture fine-grained and semantic information of pictures [23, 25, 21, 35]. Initially, the CLS embeddings of the CLIP image encoder was used to align voxels in Mind-Reader [21]. Considering the CLS embeddings alone provide limited information for fMRI data, BrainClip [23] and UniBrain [25] introduce the text annotations of images to assist in aligning fMRI voxels with image representations. Brain-diffuser [27] aligns voxel representations with the last hidden layer of CLIP using large-scale linear models ($257\times 14$M) for the first time.
@@
To tackle the trade-off between model capacity and scalability in fMRI-to-image retrieval, we propose Lite-Mind, a lightweight, efficient, and robust brain representation learning paradigm. We redesign MindEye with our elaborate Discrete Fourier Transform (DFT) Backbone, avoiding the large MLP Backbone used in MindEye. Extensive experiments show that Lite-Mind achieves 94.6% retrieval accuracy for Subject 1 on the NSD dataset, with 98.7% fewer parameters than MindEye. Lite-Mind also proves its adaptability to smaller brain datasets and establishes a new state-of-the-art zero-shot classification on the GOD dataset. Our contributions are summarized as follows:
  • We demonstrated that the Fourier Transform has the advantage of lightweight and efficient for fMRI representation learning in frequency domain.
  • We proposed a new Brain Representation Network DFT Backbone based on Discrete Fourier Transform, as an efficient and robust fMRI representation network for fMRI-to-image retrieval. Its effectiveness, efficiency, and lightweight are demonstrated.
  • We migrated Lite-Mind to various downstream tasks, and demonstrated its robustness. It achieves state-of-the-art zero-shot classification performance on the GOD dataset.
@@ 2.1 Brain Visual Decoding
The study of visual decoding using fMRI in the human brain has been a long-standing endeavor. In the most important fMRI-image reconstruction and partial fMRI-image retrieval tasks, the foundation is how to align fMRI signals to the image representation space or intermediate space. Due to the low signal-to-noise ratio of fMRI, the initial research emphasized using models based on linear regression to extract fMRI signals and images into the intermediate space for functional decoding of the human brain or reconstruction of faces and natural scenes [38]. Further research utilizes pre-trained VGG to enrich hierarchical image features [13, 36, 7]. With the development of cross-modal tasks, the image-text pre-training model CLIP has been introduced into fMRI-image research. Mind-reader [21] is the first to adopt a contrastive learning approach, using the nsdgenal ROI region of NSD data and a shallow resnet-like model to align fMRI with the CLS embeddings output by the CLIP image encoder. Mind-Reader did not achieve good performance in fMRI-image retrieval when using InfoNCE loss. BrainClip [23] uses a VAE model for retrieval tasks while still using the CLS embeddings. Brain diffuser [27] introduced a fine-grained representation of the image of the last hidden layer in CLIP. But they used 257 different linear regression models to align each layer of the 257 hidden layers, resulting in the model being too cumbersome. Mind-Vis [5] conducted self-supervised mask pre-training on fMRI signals for another dataset, BOLD5000, and demonstrated that voxel patches can be used to process fMRI voxel values. MindEye [35] uses a large MLP backbone for contrastive learning, aligning the last hidden layer $257\times 768$ of CLIP, and using diffusion prior in DALLE·2 [31] for the first time to narrow the disjointed vectors. At the same time, it creatively proposed a mixed contrast loss MixCo as a data augmentation method to make the large model converge. The recent trend in brain fMRI decoding research tends to use larger models to achieve better performance on downstream tasks [42, 16], ignoring the privacy and deployment issues discussed above. However, to the best of our knowledge, no previous research has focused on lightweight and efficient brain networks.
@@ 2.2 Fourier Transform in Deep Learning
Fourier Transform plays a vital role in the area of digital signal processing. It has been introduced to deep learning for enhanced learning performance [11, 8, 45], and has demonstrated excellent performance in many fields, such as computer vision, natural language processing, and time series analysis. In the field of computer vision, GFNet [32] utilizes Fast Fourier Transform(FFT) to convert images to the frequency domain and exchange global information between learnable filters. As a continuous global convolution independent of input resolution, Guibas et al. [12] design the Adaptive Fourier Neural Operator(AFNO) frame token mixing. Xu et al. [43] devise a learning-based frequency selection method to identify trivial frequency components and improve image classification accuracy. As for natural language processing, Lee-Thorp et al. [20] use the Fourier Transform as a text token mixing mechanism on text classification. In research of time series forecast, Fourier Transform has deeper applications  [3, 18, 17, 44, 46, 19]. To increase the accuracy of multivariate time-series forecasting, Cao et al. [3] propose a spectral temporal graph neural network (StemGNN), which mines the correlations and time dependencies between sequences in the frequency domain. Yang et al. [44] propose bilinear temporal spectral fusion (BTSF), which updates the feature representation in a fused manner by explicitly encoding time-frequency pairs and using two aggregation modules: spectrum-to-time and time-to-spectrum. Yi et al. [47] prove that frequency-domain MLP is more efficient than time-domain MLP. Processing fMRI frequency domain information is more helpful for understanding the mechanism of collaborative operation within the human brain. However, there is currently no research on processing fMRI signals in the frequency domain for representation learning.
Figure 2: Overview of our Lite-Mind. Figure (a) is the architecture of the MLP Backbone of MindEye. Figure (b) is the architecture of the DFT backbone and Retrieval Pipeline. fMRI voxels are inputted into DFT Backbone to obtain voxel embeddings for downstream tasks.
@@ 3.1 Overview
As illustrated in Figure 2, Lite-Mind comprises two main components: 1) DFT Backbone consists of fMRI Spectrum Compression and fMRI Token Decoupling in Frequency-domain, mapping flattened voxels to an intermediate space. 2) Retrieval pipeline contains Diffusion Projector, Contrastive Learning, and Downstream tasks with the voxel embeddings from the DFT backbone.
@@ 3.2 DFT Backbone
Sampling an fMRI-image pair ($x$,$y$) from dataset $D$, $x$ denotes the fMRI data and $y$ denotes the paired image.
3.2.1 fMRI Spectrum Compression
Patchify and Embedding The fMRI data $x$ is one-dimensional data composed of voxels. we firstly divide $x$ into $n$ non-overlapping patches $x=[x_{1},x_{2},...,x_{n}]$ using zero-padding, and adopt convolutions with positional encoding to obtain multiple tokens $t=[t_{1},t_{2},...,t_{n}]$.
Filter Block. The spectrum of voxel tokens is obtained by 1D DFT as follows:
$X[k]=F(t)=\sum_{i=1}^{n}t_{i}e^{-ki(2\pi/n)j}$ (1)
Where $X\in\mathbb{C}^{n\times d}$ is a complex tensor, $X[k]$ is the spectrum of $t$ at the frequency $2\pi k/n$, $F(\cdot)$ denotes the 1D DFT along the voxel token dimension. $i$ denotes the $i$-th token, and $j$ denotes unit imaginary.
After obtaining the spectrum of voxel tokens, we introduce learnable filters for filtering processing. Voxel spatial features are effectively consolidated within each frequency element, enabling the extraction of informative features from voxels through the point-wise product in the frequency domain. We use $\textbf{K}=[\textbf{k}_{1},\textbf{k}_{2},...,\textbf{k}_{M}]$ to represent the filter library, where $M$ is the number of filters in the filter library. Spectrum $\hat{X}$ after filtering and compression is obtained as follows:
$\hat{X}=\sum_{m=1}^{M}\frac{1}{n}{|X|}^{2}\odot\textbf{k}_{m}cos(\frac{(2m-1)\pi}{2M})$ (2)
where $\odot$ is the element-wise multiplication, $|X|^{2}$ is the power spectrum of $X$. The $|X|^{2}$ operation smooths the spectrum, highlighting the main components of the fMRI spectrum from a global perspective. It also facilitates voxel information concentration and noise filtering. $cos((2m-1)\pi/2M)$ adopts the filterling form of Discrete Cosine Transform (DCT), which compacts better energy and can aggregate the more important information in fMRI voxels. Its combination of application with the channel filter library K allows for efficient frequency domain feature compression and noise denoising for fMRI.
Finally, we employ Inverse Discrete Fourier Transform (IDFT), denoted by $F^{-1}(\cdot)$, to convert the spectrum back into the spatial domain:
$\hat{t}\leftarrow F^{-1}(\hat{X})$ (3)
3.2.2 fMRI Token Decoupling in Frequency-domain
To align voxel tokens to image tokens, we decouple the filtered and compressed fMRI tokens again in frequency domain. Similarly, we follow Equation (2) to convert $\hat{t}$ to the frequency domain, then decouple tokens by a MLP-like model in frequency domain(FreMLP).
FreMLP. Following Equation (2), the complex number output of $\hat{t}$ after DFT is $\hat{X}\in\mathbb{C}^{n\times d}$. Given a complex number weight matrix $\mathcal{W}\in\mathbb{C}^{n\times n^{\prime}}$, and a complex number bias $\mathcal{B}\in\mathbb{C}^{n^{\prime}}$, then the frequency-domain MLPs for token decoupling can be formulated as:
$X^{\prime}={\sigma(\hat{X}^{T}\mathcal{W}+\mathcal{B})}^{T}$ (4)
Where $X^{\prime}\in\mathbb{C}^{n^{\prime}\times d}$ is the final output, and $\sigma$ is the activation function. As both $\hat{X}$ and $\mathcal{W}$ are complex numbers, according to the rule of multiplication of complex numbers(see Appendix D.1 for details), we further extend the Equation (5) to:
$\begin{split}X^{\prime}&=({\sigma(Re(\hat{X}^{T})\mathcal{W}_{r}-Im(\hat{X}^{T})\mathcal{W}_{i}+\mathcal{B}_{r})}\\ &+j\sigma(Re(\hat{X}^{T})\mathcal{W}_{i}+Im(\hat{X}^{T})\mathcal{W}_{r}+\mathcal{B}_{i}))^{T}\end{split}$ (5)
Where $Re(\cdot)$ denotes the real part of $\hat{X}^{T}$, $Im(\cdot)$ denotes the imaginary part, $\mathcal{W}=\mathcal{W}_{r}+j\mathcal{W}_{i}$ and $\mathcal{B}=\mathcal{B}_{r}+j\mathcal{B}_{i}$. Due to the more significant features of filtered and compressed tokens in the frequency domain, we replace the final FC layer of MindEye’s MLP Backbone with FreMLP. For more theoretical proof and implementation of FreMLP, please refer to the Appendix D.2. Finally, we follow IDFT $t^{\prime}\leftarrow F^{-1}(X^{\prime})$ to get final tokens $t^{\prime}$, and use voxel embedding $f$ to represent them.
3.3 Retrieval Pipeline
3.3.1 Optimization Objective
The optimization objective for the retrieval task is defined as follows:
@@ 3.3.1 Optimization Objective
where $\omega$ is the weight of our DFT backbone, $SIM(\cdot)$ denotes the cosine similarity, and $CLIP(\cdot)$ denotes using CLIP model to extract image embedding.
3.3.2 Diffusion Projector
When searching on a large-scale dataset like LAION-5B, image-to-image performs better than fMRI-to-image and can find more similar images. Therefore, we used a diffusion projector to translate the disjointed voxel CLS embeddings of the DFT Backbone output, using a DALLE · 2[31] similar diffusion model to obtain the image CLS embeddings and retrieve it on LAION-5B: $\mathcal{V^{\prime}}=Diffusion(f)$. The Diffusion Projector consists of a transformer for image CLS embedding forecasting.
3.3.3 Contrastive Learning
Due to the inherent differences between different modalities, we use CLIP’s contrastive loss to train the DFT backbone. Sampling a batch $B$ from voxel-image pairs, the contrastive loss is defined below:
$L_{contr}=-\frac{1}{|B|}\sum_{s=1}^{|B|}\log\frac{\exp(f_{s}^{\top}\cdot\mathcal{V}_{s}/{\tau})}{\sum_{i=1}^{|B|}\exp({f_{s}^{\top}\cdot\mathcal{V}_{i}}/{\tau})},$ (7)
where $f_{s}$ is the $s$-th voxel embeddings, $\mathcal{V}_{s}$ is its corresponding image embeddings, $\tau$ is a temperature factor.
To perform image-to-image retrieval on LAION-5B, we use MSE loss to constrain the generation of approximate image embeddings:
$L_{mse}=\frac{1}{|B|}\sum_{s=1}^{|B|}{||\mathcal{V}_{s}-\mathcal{V^{\prime}}_{s}||}^{2}_{2}$ (8)
Accordingly, the training loss is defined below:
$L=L_{contr}+\alpha L_{mse}$ (9)
In general, when $\alpha=0$, Lite-Mind is trained to output voxel embeddings for retrieval on the NSD test set. However, for LAION-5B retrieval, $\alpha$ is non-zero to map the voxel CLS embeddings outputted by Lite-Mind to image CLS embeddings, enabling online retrieval.
3.3.4 Downstream tasks